math homework: carrie is three times as old as david. in five years carrie will be only twice as old as david. how old will Carrie be in five years
Found 2 solutions by ikleyn, lwsshak3:
Answer by ikleyn(52790)


Let  x  be the David' present age.
Then Carrie's present age is  3x.

In  5  years Carries age will be  3x+5,  and David's age will be  x+5.

According to the condition,  the equation holds

3x + 5 = 2(x+5).

Solve it:

3x + 5 = 2x + 10,

3x - 2x = 10 - 5,

x = 5.

So,  David is 5 years old now.  Carrie is 3*5 = 15 years old now.


Answer by lwsshak3(11628)

math homework: carrie is three times as old as david. in five years carrie will be only twice as old as david. how old will Carrie be in five years
let x=David's age
3x=Carrie's age
In 5 yrs:
x+5=david's age
3x+5=Carrie's age
3x+5=2(x+5
3x+5=2x+10
x=5
3x+5=20
. how old will Carrie be in five years? 20